OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (anonymous):

The x and y intercepts are always going to be when the other is 0. Try plugging 0 in for each variable (one at a time) and your result should be the coordinate.

OpenStudy (anonymous):

can you give some example please

OpenStudy (anonymous):

(3/2 , -3) (0 , -5 ) (1/2 , -7)

jimthompson5910 (jim_thompson5910):

y-intercept y = -2x - 3 y = -2(0) - 3 y = ???

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

so the y-intercept is (0, -3)

OpenStudy (anonymous):

x- intercept y intercept (0, 3/2 ) (0, -3) (1 , 0) (1, -5) ( 2, 1/2 ) (2, -7)

OpenStudy (anonymous):

now is it correct

jimthompson5910 (jim_thompson5910):

the x-intercept is found by plugging in y = 0 and solving for x y = -2x - 3 0 = -2x - 3 3 = -2x 3/(-2) = x x = -3/2 or x = -1.5 so the x-intercept is the point (-3/2, 0) or (-1.5, 0)

OpenStudy (anonymous):

but i thought when we will -2 on the let side it will become +ve

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

when we moved -2 from the right side to the left side will it not become +2

jimthompson5910 (jim_thompson5910):

no you move the -3 over first

jimthompson5910 (jim_thompson5910):

then you move the -2 over

jimthompson5910 (jim_thompson5910):

so you first add 3 to both sides

jimthompson5910 (jim_thompson5910):

then you divide both sides by -2

OpenStudy (anonymous):

so the coordinates i have given are wrong

jimthompson5910 (jim_thompson5910):

for the x-intercept, yes

jimthompson5910 (jim_thompson5910):

it should be (-3/2, 0) or (-1.5, 0)

OpenStudy (anonymous):

so we just have to give 1 coordinate

jimthompson5910 (jim_thompson5910):

you mean one point? yes you give one point because there is only one x-intercept

jimthompson5910 (jim_thompson5910):

and there is only one y-intercept

OpenStudy (anonymous):

i means we will not be imagining 1 instead of zero now

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

for example let x= 1 instead of zero

jimthompson5910 (jim_thompson5910):

ah i see, so you're looking for other points now

jimthompson5910 (jim_thompson5910):

you can do that, it just won't be called the x or y intercept

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

yw